OpenStudy (anonymous):

Which expression is not equivalent to the other expressions? zero point five times four raised to the second power plus ten two point five times the quantity six minus two raised to the second power, end quantity two raised to the third power divided by ten plus four point two the quantity six plus three raised to the second power, end quantity, divided by three

OpenStudy (fibonaccichick666):

can you write them all out in symbols?

OpenStudy (anonymous):

\[0.5\times4^{2}+10\] \[2.5\times(6-2^{2})\] \[2^{3}\div10+4.2\] \[(6+3^{2})\div3\]

OpenStudy (anonymous):

I don't understand, which expressions are we comparing?

OpenStudy (anonymous):

those 4..

OpenStudy (anonymous):

Lol but "Which expression is not equivalent to the other expressions?" doesn't make sense to me here.

OpenStudy (fibonaccichick666):

now are you familiar with the order of operations?

OpenStudy (fibonaccichick666):

@assaulth3ro911 which one doesn't give you the same end result?

OpenStudy (anonymous):

Is that what they were asking...? Oh darn, now I have to do all the equations. Oh well. (x

OpenStudy (fibonaccichick666):

nah, you only need to do 3 but I have the answer ready for when @twilaswift gets there

OpenStudy (anonymous):

0.5×4^2+10 = 18 2.5×(6−2^2) = 5 2^3÷10+4.2 = 5 (6+3^2)÷3 = 5

OpenStudy (anonymous):

The answer would be A.

OpenStudy (anonymous):

Not unless you show how you got there... of which...... I didn't do, but it was pretty simple to explain, so we explained it just fine I guess. (: